OpenStudy (ayonnaleflore):

Find the zeros of the polynomial function. f(x) = x2 – 12x + 20 4, 5 2, 10 –4, –5 –2, –10

OpenStudy (mathmale):

You posted this earlier and I responded in depth. Why are y ou posting this question again? I'm not eager to repeat what I typed earlier in an effort to help you.

OpenStudy (kittiwitti1):

Just set the problem equal to 0 and solve for x ^-^

OpenStudy (xmissalycatx):

Do you know how to factor? >_>

OpenStudy (ayonnaleflore):

x^2 -12x+20=0 x^2-12x=-20 now once i get to this point what should i do

OpenStudy (kittiwitti1):

The quadratic formula also works.\[f(x)=ax^{2}+bx-c\rightarrow x=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (kittiwitti1):

Oh, you're completing the square?

OpenStudy (mathmale):

Keep this equation as given: x^2 -12x+20=0 It does not help to move that 20 to the right side.

OpenStudy (kittiwitti1):

Yes, what @mathmale said... completing the square doesn't help you find the zeros. I believe it only helps you find the vertex form of a quadratic function

OpenStudy (mathmale):

As I shared with you earlier, methods of solving this equation include: 1. graphing 2. completing the square 3. factoring 4. quadratic formula 5. factoring by grouping Please identify the approach that YOU want to use.

OpenStudy (xmissalycatx):

But this is a simple question. All you need to do is factor to find the 0's.

OpenStudy (mathmale):

Yes: that'd be the easiest approach. But learning to use other approaches would be helpful as well.

OpenStudy (xmissalycatx):

I meant to send that in reply to kitti, my apologies :P

OpenStudy (kittiwitti1):

xD Well, the quadratic formula also works as the "long way round" when you can't factor or get stuck doing that... ^-^

OpenStudy (kittiwitti1):

Oh, so you CAN use completing the squ- wait yeah I'm dumb lol Do you know how to continue solving using complete the square?

OpenStudy (mathmale):

@ayonnaleflore: Your input, please? Questions? Comments?

OpenStudy (mathmale):

Yes, completing the square would definitely "work" here. But: factoring would be the fastest method.

OpenStudy (ayonnaleflore):

no i dont know how to continue by completing the square @kittiwitti1

OpenStudy (kittiwitti1):

http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (mathmale):

Then please cross that off the list. Which other method would you like to use?

OpenStudy (kittiwitti1):

The link I posted helps with completing the square, but you can also use other options, as we have said before ☻

OpenStudy (ayonnaleflore):

factoring by grouping but im not sure how to do that

OpenStudy (mathmale):

As I shared with you earlier, methods of solving this equation include: 1. graphing 2. completing the square 3. factoring 4. quadratic formula 5. factoring by grouping Please identify the approach that YOU want to use.

OpenStudy (ayonnaleflore):

so how would i start it off???

OpenStudy (kittiwitti1):

I don't think you can factor by grouping when there's a trinomial (three terms). I thought it required four terms.. :\

OpenStudy (ayonnaleflore):

wait doesn't it need four terms to factor by grouping

OpenStudy (kittiwitti1):

lol yeah, my question exactly xD

OpenStudy (ayonnaleflore):

ok so how about factoring??

OpenStudy (mathmale):

You can "make" four terms during factoring by grouping. Separate -12x into two parts that combine to produce -12x. Factor by grouping. Or, if you prefer, just factor. How would you do that? Hint: What are possible factors of 20?

OpenStudy (kittiwitti1):



OpenStudy (kittiwitti1):

Oh right. I suppose you could do that

OpenStudy (mathmale):